### v2

# supplement.pdf

Abstract
We provide a framework to determine the upper bound for the complexity of computing the exact Heisenberg representation of a given operator with respect to a Hamiltonian. Working in Heisenberg picture, we show that each Hamiltonian defines an equivalence relation, causing the operator space to be partitioned into equivalence classes. Any operator within a specific class never leaves it during the evolution. We provide a method to determine the dimension of the equivalence classes and evaluate it for various models, such as the $XY$ chain and Kitaev model on trees. For classes with a dimensionality that is not excessively large, one can use our framework to compute exact Heisenberg representation of any operator within this class. Our findings reveal that the complexity of operator evolution in these models grows from the edge to the bulk, which is physically manifested as suppressed relaxation of qubits near the boundary. Our methods are used to reveal several new cases of simulable quantum dynamics, such as $XY$ evolution augmented with certain non-Clifford single-qubit gates. We also introduce a $XY$-$ZZ$ model where the dimension of an equivalence class, while exponential in the system size, grows considerably slower than could be naively anticipated, which entails practical computational advantage when simulating finite systems. Furthermore, we demonstrate how to apply our method to time-dependent Hamiltonians and dissipative systems.
Introduction
Solving time dynamics of quantum many-body problems is typically an extremely difficult task. In a brute force computation of such systems, one requires solving a coupled differential equations with an exponential number of variables with respect to the system size. In the Heisenberg formulation, an initially local observable evolves into an exponentially complex sum of non-local operators [1, 2, 3, 4, 5, 6]. This proliferation of operators in the Heisenberg picture, called operator growth, has been explored recently in numerous closed [4, 6, 1, 2, 7, 8, 9, 10, 11] and open quantum systems [12, 13, 14, 15, 16, 17, 18]. Understanding operator complexity is crucial for determining the boundary between solvable and unsolvable problems in quantum dynamics, distinguishing between reachable and non-reachable quantum states, and understanding how information propagates through quantum systems. In a quantum computing context, the simulability of quantum circuits is a crucial question in the context of quantum advantage and other applications [19, 20, 21, 22, 23, 24, 25].
In some special cases, the space where the operator evolution occurs may not necessarily cover the full operator space. For instance, in free-fermionic spin chains, the $z$-spin projection operator under Heisenberg evolution only scales polynomially with the system size [26, 27]. This restricted evolution explains the solvability of various dynamic [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] and transport [41, 42, 43, 41, 44, 45] problems, as well as the classical simulability of matchgate circuits [26, 27, 46, 47, 48]. Moreover, different local operators within the same system can evolve in non-overlapping subspaces with varying dimensions and growth patterns. Examples include the $z$- and $x$-spin projections in the $XX$ model [2], or edge modes decoupled from the bulk [49, 50, 51, 52] discovered in various models. The dimensionality of this subspace, which we call the Operator Evolution Dimension (OED), provides a natural upper bound for the complexity of simulating the corresponding quantum dynamics problem.
In this paper, we introduce a framework to classify the OEDs of different local operators, which allows for a powerful way of understanding of the difference between simulable and non-simulable quantum dynamics problems. Given a spin-$1/2$ Hamiltonian, we show that one may define a equivalence relation on the set of Pauli strings, such that the span of each equivalence class is closed with respect to commutation with the Hamiltonian. We then study how specific Hamiltonians segregate the space of Pauli strings into these equivalence classes and compute the corresponding OEDs for operators within these classes. In the case of $XY$ dynamics we derive exact expressions for OEDs as integer-valued polynomials. Such knowledge of OEDs gives insight into the complexity of quantum dynamics of the system, which we illustrate on several models, including some which are non-free fermion models. In particular, we fully characterize OEDs of all local operators in $XY$ Hamiltonians. To our knowledge, such a characterization was only available before for a very limited set of local operators that are quadratic in fermionic representation.
Equivalence classes of Pauli strings
In the operator space of $L$ spin-1/2 particles, we define a Pauli string as $P=S_{1}\otimes\dots\otimes S_{L}$, where $S_{i}\in\{I_{i},X_{i},Y_{i},Z_{i}\}$. Here $X_{i},Y_{i},Z_{i}$ are Pauli matrices acting on $i$th site and $I_{i}$ is the identity matrix. There are $4^{L}$ different Pauli strings in total that constitute a complete oeprator basis $\mathcal{P}=\{P_{n}\}^{4^{L}}_{n=1}$.
Consider an operator $A\in\mathcal{P}$, such that $A\equiv P_{n}$. The aim in a time dynamical problem is to find the Heisenberg representation of this operator $A(t)$, given by following expansion:
$A(t)\equiv\sum\limits^{D}_{m=1}f_{m}(t)P_{m},$ (1)
where $f_{m}(t)$ are time-dependent functions that should be obtained from the solutions of Heisenberg equations and $D$ is the OED of the operator $A$. The evolution is governed by the Hamiltonian $H$ that can be represented as a sum of $M$ Pauli strings,
$H=\sum\limits_{n=1}^{M}h_{n}H_{n},$ (2)
where $H_{n}\in\mathcal{P}$ are Pauli strings (referred to as Hamiltonian strings) and $h_{n}$ are real numbers.
Given the set of Hamiltonian strings =$\mathcal{H}=\{H_{n}\}_{n=1}^{M}$, we can define an equivalence relation among Pauli strings (denoted by $\sim$) within the set $\mathcal{P}$. Specifically, we consider strings $P_{n}$ and $P_{m}$ as equivalent if they coincide or if one can be derived from another after $Q$ rounds of commutation with Hamiltonian strings from $\mathcal{H}$. More specifically, there exists $\{H_{k_{1}},\dots,H_{k_{Q}}\}\in\mathcal{H}$ and a complex number $a$ such that:
$P_{n}=a[H_{k_{Q}},\dots,[H_{k_{2}},[H_{k_{1}},P_{m}]]\dots].$ (3)
This binary relation satisfies the axioms of equivalence relations. This implies reflexivity $P_{n}\sim P_{n}$, symmetry [^1] This follows from $[H_{k},[H_{k},P]]=4P$ for any two non-commuting Pauli strings $H_{k}$ and $P$. $P_{n}\sim P_{m}\iff P_{m}\sim P_{n}$, and transitivity: if $P_{n}\sim P_{m}$ and $P_{m}\sim P_{l}$ then $P_{n}\sim P_{l}$.
[…]
In general, for a given Hamiltonian, the equivalence class of any operator can be determined by algorithmically generating all strings that result from commutation with the Hamiltonian strings. The amplitudes $f_{m}(t)$ in the Heisenberg representation can be computed by solving the Heisenberg equations within an isolated class. In the case of non-integrable dynamics at late times, these amplitudes follow a Porter-Thomas distribution, and deviate from it in the integrable case, see Supplementary material for details [53].
Figure 1: Schematic representation of the partitioning of the operator space of Pauli strings into $12L+$ 1 equivalence classes by the $XY$ Hamiltonian (Disordered $XY$-spin chains.). Each segment is labeled with a representative observable from the corresponding class. Operators that belong to a certain class never leave it during the evolution. Pauli strings that are integrals of motion are depicted as dots at each end of the ellipse.
Disordered $XY$-spin chains.
Let us consider a simple application of the equivalence class framework by examining the disordered $XY$-spin chain, with Hamiltonian
$H_{\text{XY}}$ $=\sum\limits^{L}_{i=1}J^{xx}_{i}X_{i}X_{i+1}+J^{yy}_{i}Y_{i}Y_{i+1}+$
$J^{xy}_{i}X_{i}Y_{i+1}+J^{yx}_{i}Y_{i}X_{i+1}+h^{z}_{i}Z_{i},$ (4)
Here the $J$ and $h$ coefficients take arbitrary real values. Unless stated otherwise, we will assume open boundary conditions.
We start with the well-known classes $\mathcal{A}^{1}_{\text{XY}}$ and $\mathcal{A}^{2}_{\text{XY}}$, which can be generated from the strings $X_{1}$ and $Z_{i}$, respectively. The class $\mathcal{A}^{1}_{\text{XY}}$ is known as Majorana strings and has the following elements:
$\mathcal{A}^{1}_{\text{XY}}=$$\left\{X_{1},Z_{1}X_{2},Z_{1}Z_{2}X_{3},\dots,Z_{1}\dots Z_{L-1}X_{L},\right.$
$\left.\>\;Y_{1},\,Z_{1}Y_{2},\,Z_{1}Z_{2}Y_{3},\;\dots,Z_{1}\dots Z_{L-1}Y_{L}\right\},$ (5)
which has dimension $D^{1}_{\text{XY}}=2L$. The class $\mathcal{A}^{2}_{\text{XY}}$ consists of special Pauli strings known as Onsager strings [54, 40]. An Onsager string is either a single Pauli matrix $Z_{j}$ or a product of Pauli matrices on consecutive sites with matrices $X$ or $Y$ at the ends and matrices $Z$ in the middle, e.g. $Y_{3}Z_{4}Z_{5}Z_{6}X_{7}$. There are $D^{2}_{\text{XY}}=2L^{2}-L$ Onsager strings. Both these classes have an OED that are low-order polynomials in $L$. This can be considered the fundamental reason that particular observables in these classes can be simulated efficiently.
Interestingly, these are not the only classes that are present for the $XY$ Hamiltonian. In fact we find that $\mathcal{P}$ is divided into $K=2L+1$ equivalence classes as $\mathcal{P}=\mathcal{A}^{0}_{\text{XY}}\bigcup\dots\bigcup\mathcal{A}^{2L}_{\text{XY}}$. The full set of classes can be denoted by $\mathcal{A}^{2n-1}_{\text{XY}}\equiv\mathcal{A}_{\text{XY}}[X_{n}]$ and $\mathcal{A}^{2n}_{\text{XY}}\equiv\mathcal{A}_{\text{XY}}[\prod^{n}_{m=1}Z_{m}]$, where $n\in[1,L]$. In addition to these there is also the trivial identity class $\mathcal{A}^{0}_{\text{XY}}=\{I_{1}\otimes\dots\otimes I_{L}\}$ which is always present in any model. The partition of $\mathcal{P}$ into equivalence classes induced by the Hamiltonian (Disordered $XY$-spin chains.) is graphically represented in Fig. 1.
Figure 2: The growth of OEDs for various operators as a function of the system size $L$ and lattice index $i$. Log-log plot illustrating the (a) OEDs of different operators as labeled in the $XY$ chain (Disordered $XY$-spin chains.); (b) OEDs of operators $X_{i}$ in the $XY$ chain $\mathcal{A}_{\text{XY}}[X_{i}]$ as a function of lattice index $i$; (c) sum of OEDs for single site operators in the Kitaev chain $D^{i}_{\text{Kit-loc}}=D_{\text{Kit}}[X_{i}]+D_{\text{Kit}}[Y_{i}]+D_{\text{Kit}}[Z_{i}]$. (d) Semi-log plot showing numerically obtained values of OEDs for $Z_{1}$ operator in the $XY$-$ZZ$ chain and the corresponding exponential fit $\sim 11.03e^{0.365413L}$.
Let us now determine the dimension of these equivalence classes (see also Supplementary Material [53]). For an arbitrary system size $L$, the OED of the equivalence class $\mathcal{A}^{N}_{\text{XY}}$ can be written as a polynomial of degree $N$:
$D^{N}_{\text{XY}}(L)=\sum\limits^{N}_{j=0}k^{N}_{j}L^{j}.$ (6)
The coefficients $k^{N}_{0},\dots,k^{N}_{N}$ for $N\leq L$ are determined as solutions of the following system of equations:
$V(0,\dots,N)\vec{K}^{N}=\vec{D}^{N}_{\text{XY}}.$ (7)
Here $V(0,1,\dots,N)$ is a Vandermonde matrix, $\vec{K}^{N}\equiv(k^{N}_{0},\dots,k^{N}_{N})$, and the vector of values $\vec{D}^{N}_{\text{XY}}=(D^{N}_{\text{XY}}(0),\dots,D^{N}_{\text{XY}}(N))$ is given by
$D^{N}_{\text{XY}}(m)=\begin{cases}0&\;m\leq\lceil N/2-1\rceil\\ D^{\lceil 2(m-N/2)\rceil}_{\text{XY}}(m)&\;m\geq\lceil N/2\rceil\\ 4^{N}-2\sum\limits^{N-1}_{n=0}D^{n}_{\text{XY}}(m)&\;m=N\end{cases}.$ (8)
For a specific $N$, the vector of values is determined from the previous $N-1$ polynomials. Specifically, $D^{N}_{\text{XY}}(m)$ is equal to zero for the first several values of $m$, as on a lattice of length $m$ there is simply no equivalence class that corresponds to $N$. The remaining values are determined from the relation $D^{n}_{\text{XY}}=D^{2L-n}_{\text{XY}}$. The same relation provides the polynomial in the case $N\geq L+1$. The last line in (8) is derived from the fact that the sum of dimensions of all classes must be equal to $4^{L}$. Therefore, starting from classes $\mathcal{A}^{1}_{\text{XY}}$ and $\mathcal{A}^{2}_{\text{XY}}$ one can iteratively construct all the other classes and compute their dimensionalities. Despite the extensive research history of some of these models, to our knowledge the complexity of arbitrary local operators has not been discussed.
For example, the expression for $D^{5}_{\text{XY}}$ is found to be:
$D^{5}_{\text{XY}}=\frac{4}{10}L-\frac{5}{3}L^{2}+\frac{7}{3}L^{3}-\frac{4}{3}L^{4}+\frac{4}{15}L^{5}.$ (9)
The above integer-valued polynomial gives the OED of the operators $X_{3}$ and $Y_{3}$. This means that determining the exact dynamics of these operators in the basis of Pauli strings would require solving $D^{5}_{\text{XY}}$ linear equations. Figure 2(a) shows some other examples on a logarithmic plot, showing the polynomial growth of these equivalence classes. For an arbitrary lattice index $X_{n}$, the dimensionality of the corresponding class is given by a polynomial of degree $12n-$ 1. In a computational complexity sense, this makes calculating the dynamics of any arbitrary Pauli string in the $XY$ chain efficient, in the sense that the OED is polynomial. However, towards the middle of the chain the degree of the polynomial may be prohibitively large such that practically larger chains are inaccessible. In Fig. 2(b), we plot the dependence of $D_{\text{XY}}[X_{i}]$ on the lattice index $i$, showing a peak in the middle of the chain and smaller values near the edges. This example demonstrates the edge-to-bulk growth of OED within the $XY$-chain. This implies the greater computational overhead to study dynamics of observables in the bulk of the system.
Figure 3: Single-qubit relaxation dynamics for the initial state $\rho_{\text{init}}=|\psi_{i}\rangle\langle\psi_{i}|\otimes I^{\otimes L-1}/2^{L-1}$, here $|\psi_{i}\rangle$ is the state of $i$-th qubit, the rest of the system is at the infinite temperature. (a) $XY$-chain (Disordered $XY$-spin chains.), $J^{\alpha\beta}_{i}=1$, $h^{z}_{i}=10$, $L=12$, $|\psi_{i}\rangle=|+\rangle_{i}$ (b) 1D Kitaev chain (10), $J^{x}_{i}=J^{y}_{i}=J^{z}_{i}=1$, $L=30$, $|\psi_{1}\rangle=|1\rangle_{1}$, $|\psi_{2}\rangle=|+\rangle_{2}$, $|\psi_{3}\rangle=|-\rangle_{3}$.
To show that the increase in complexity from the edge to the bulk has also physical ramifications, we compute the relaxation dynamics of a single qubit at infinite temperature. The $XY$ chain is known to have boundary modes that are fully decoupled from the bulk excitations [49, 52]. As such we expect qubits on the edges to maintain coherence to a better extent that those in the bulk. We consider the $XY$ Hamiltonian with the uniform couplings $J^{\alpha\beta}_{i}=1$ and a large magnetic field $h^{z}_{i}=10$, to induce fast oscillations in the $XY$-plane so that we may observe the coherence decay [55, 56]. As the initial state, we consider $\rho_{\text{init}}\propto|+\rangle_{i}\langle+|_{i}\otimes I^{\otimes L-1}$, that corresponds to $i$th qubit being fully polarized along $x$-axis and the rest of the system being at the infinite temperature state. Although the system under consideration is integrable, a single qubit placed inside an infinite temperature bath still rapidly relaxes to its equilibrium value. This, however, does not occur for the edge qubit, as its evolution takes place in a small subspace that is decoupled from the bulk. This is manifested by the collapse and revival it undergoes during evolution, as seen in Fig. 3(a.1). We observe that qubits adjacent to the edge qubit also exhibit collapse and revival, although they are less pronounced, with the amplitude vanishing as one moves towards the bulk, see Fig. 3(a.2)(a.3). The revivals can be explained by the fact that the OED of the corresponding operator near the edge is still relatively small, so the initial qubit polarization cannot dissipate swiftly due to the limited number of degrees of freedom available for dissipation.
Mixing of equivalence classes by perturbations and quenches.
Introducing perturbations to a Hamiltonian affects the equivalence classes of the model. For example, applying periodic boundary conditions to the Hamiltonian (Disordered $XY$-spin chains.) changes the equivalence classes such that instead of having a unique class for each operator $X_{n}$, only two classes distinguished by odd and even lattice indices are present, each exhibiting exponential dimensionality $4^{L-1}$. The class $\mathcal{A}^{2}_{\text{XY}}$ remains composed of Onsager strings, but since the lattice is periodic its dimensionality doubles, resulting in $D^{2}_{\text{XY-PBC}}=2D^{2}_{\text{XY}}$. This example illustrates how the same perturbation can affect OEDs of different operators in a drastically different way.
Figure 4: (a) Growth of OED $D_{\text{XY}}[Z_{1}]$ of the $Z_{1}$ operator during the SWAP gate perturbed $XY$-evolution. Blocks $U_{\text{XY}}=e^{-itH_{XY}}$ describe evolution governed by the Hamiltonian (Disordered $XY$-spin chains.). As nearest-neighbour SWAP gates are applied, $D_{\text{XY}}[Z_{1}]$ increases as $O(L^{2+2m})$, where $m$ is the number of applied gates. (b) Illustration of a Kitaev Hamiltonian (10) on the tree. (c) Geometry of 1D Kitaev chains with and without loops.
Another example of a perturbation is applying SWAP-quenches, when a $XY$ Hamiltonian evolution is interrupted by a nearest-neighbour SWAP-gate, see Fig. 4(a). Examining the equivalence classes of $Z_{i}$, it is found that a single $\text{SWAP}_{j,j+1}$-gate application merges classes $\mathcal{A}^{2}_{\text{XY}}$ and $\mathcal{A}^{4}_{\text{XY}}$, into a unified class with the dimension being fourth order in $L$. Generally, applying multiple SWAP-quenches during $XY$ evolution increases the OED of $Z_{i}$ as polynomial of degree $2+2m$, with $m$ representing the number of quenches. This is consistent with the asymptotic results recently reported in [57]. Ultimately, SWAP-quenches will intermix all equivalence classes into one, achieving a dimensionality of $14^{L}-$ 1. This is consistent with the fact that $XY$ dynamics combined with SWAP-quenches is universal [46].
Not all quenches, however, merge equivalence classes. One example of this are parameter quenches in (2) $h_{n}\rightarrow h^{\prime}_{n}$ do not affect worst case OED cf. [58]. Due to this fact our formalism is directly extendable to time-dependent Hamiltonians [53]. Another example are Pauli gates $W=X_{i},Y_{i},Z_{i}$ which act on any Pauli string as $W^{\dagger}P_{n}W=\pm P_{n}$, leaving any equivalence class of Pauli strings unaffected. In case of the class $\mathcal{A}^{2}_{\text{XY}}$, it is easy to check that the phase gate $W=\sqrt{Z_{i}}$ and the gate $W=X_{i}\cos\alpha+Y_{i}\sin\alpha$, where $\alpha\in\mathbb{R}$, are also equivalence preserving quenches. Therefore $XY$ dynamics can be augmented with an arbitrary number of the above quenches without complicating its simulability. This example illustrates non-maximality of $XY$ dynamics with respect to addition of parity-violating unitaries[59].
Other Hamiltonians.
To show that our methods are applicable in a more general context, we consider other Hamiltonians. One example is the Kitaev Hamiltonian on a tree graph $G^{L}_{\text{tree}}$ with $L$ vertexes and at most three edges incident to each vertex (Fig. 4(b)). The Hamiltonian has the form [60]
$H_{\text{Kit}}=\sum\limits_{\langle i,j\rangle\in G}J^{x}_{ij}X_{i}X_{j}+J^{y}_{ij}Y_{i}Y_{j}+J^{z}_{ij}Z_{i}Z_{j},$ (10)
where for a given pair $\langle i,j\rangle$ of neighbouring vertexes two of three coupling constants $J^{x,y,z}_{ij}$ must be zero, but the remaining is non-zero. In addition, the Hamiltonian must have a proper 3-edge coloring of the graph, where each “color” represents the interaction type $XX,YY$ or $ZZ$, and proper coloring implies that no vertex belongs to two edges with the same color. For this Hamiltonian, each local operator $X_{n}$, $Y_{n}$ or $Z_{n}$ belongs to its unique equivalence class and has OEDs that can be given as at most $n$th degree polynomial of $L$. In Fig. 2(c) the growth of the OED for local Pauli operators is shown for the one-dimensional Kitaev chain (Fig. 4(c)), revealing a polynomial growth of $D^{n}_{\text{Kit-loc}}$. We find that OEDs of local operators in the Hamiltonian (10) can always be expressed as an integer-valued polynomial as long as the tree structure of the Hamiltonian is maintained, with the specific geometry influencing the form of the polynomial.
Kitaev Hamiltonians such as (10) offer a rich playground for exploring equivalence classes and the dynamics of operator growth. Such Hamiltonians contain numerous local operators that belong to equivalence classes of relatively low dimensions. Introducing external magnetic fields or loops in the graph one can merge some of these classes in a controllable manner. For example, in Fig. 2(c) we demonstrate how the total OEDs of local operators $X_{4},Y_{4},Z_{4}$ increases when the tree structure of the graph in the Hamiltonian (10) is disrupted by adding loops while maintaining its length (Fig. 4(c)). This results in a dimensional jump of $D^{n}_{\text{Kit-loc}}$ by a factor of $2^{m}D^{n}_{\text{Kit-loc}}$, where $m$ denotes the number of added loops. This relation holds within range of $n$ and $m$, as long as $2^{m}D^{n}_{\text{Kit-loc}}\ll 4^{L}$.
Finally, we introduce the $XY-ZZ$ model which again induces a nontrivial division of equivalence classes. The Hamiltonian of $XY-ZZ$ model reads:
$H_{\text{XY-ZZ}}=\sum\limits^{L}_{i=1}J^{x}_{2i-1}$ $X_{2i-1}X_{2i}+J^{y}_{2i-1}Y_{2i-1}Y_{2i}$
$+J^{z}_{2i}Z_{2i}Z_{2i+1},$ (11)
here again the coefficients $J$ are real. In case of this Hamiltonian the operators $X_{i}$ and $Y_{i}$ each belong to distinct equivalence classes with dimensionalities $D_{\text{XY-ZZ}}[X_{i},Y_{i}]=4^{L-3}$. Pairs of operators $Z_{2i-1}$ and $Z_{2i}$ for $i>1$, belong to equivalence classes with slightly bigger dimensionality scaling as $O(4^{L-3})$. While the OED of these classes are exponentially growing, some grow at a much slower rate $2^{L}$. The most intriguing case involves the edge operator $Z_{1}$, whose equivalence class dimensionality grows approximately as $D_{\text{XY-ZZ}}[Z_{1}]\simeq 11e^{0.365L}$, see Fig 2(d). The factor $0.365$ in the exponent enables a significant practical reduction in problem complexity, compared to exact diagonalization.
Dissipative dynamics.
Lastly, we also mention that the partitioning of the operator space to equivalence classes also takes place for dissipative dynamics that is described by GKSL equation [61]. Furthermore, in certain cases, dissipation also does not violate the structure of certain equivalence classes or violates it insignificantly. This is the case e.g. for the Hamiltonian (Disordered $XY$-spin chains.) with dephasing [41, 42, 41, 62, 63, 64, 44]. Thus, our methods can be directly applied to treat various (semi-)integrable dissipative models [65, 43, 40, 45] as well as for finding new ones. See supplementary material for details [53].
Discussion.
We have developed a formalism of equivalence classes which allows one to obtain the dimension of the operator evolution in the Heisenberg picture for spin-$1/2$ Hamiltonians. The approach gives a straightforward way of understanding the nature of why particular models can be time integrable, and provides a powerful tool for identifying when quantum dynamics are simulable. To illustrate our approach we examined well-studied models such as the $XY$ and Kitaev chains and recovered the known result that $X_{i},Y_{i}$ at the chain ends and all $Z_{i}$ can be efficiently computed. Despite the long history of these models, to our knowledge, the remaining equivalence classes have not been identified to date. We showed that the $X_{i},Y_{i}$ observables increase in complexity moving from edge to bulk. We also considered various perturbations that may or may not affect the OEDs of certain operators. We showed that applications of certain single-qubit gates during the evolution do not affect OED of some operators under simulation. This is interesting from a quantum circuit point of view as it implies that matchgate circuits can be augmented with such gates without penalty regarding the simulation. We examined universality-enabling gates for $XY$ evolution, specifically SWAP gates, which increases the degree of the OED linearly with the number of SWAP gates. Similarly, loops on Kitaev model on trees increases the complexity by an exponential factor in the number of loops.
To illustrate our results we mainly showed results on free-fermionic models, but they may also be applied in other contexts. We introduced the non-free fermion $XY$-$ZZ$ model, we showed that all equivalence classes have exponential dimensionality. In this model, almost each local operator still resides within its own equivalence class. Moreover, the $z$-projections of edge qubits also have OEDs that scale more slowly than those of bulk qubits. Solving the dynamics within the corresponding equivalence class will still significantly outperform brute force approach that disregard the splitting of the operator space, such as exact diagonalization. While we focused on the basis of Pauli strings for simplicity in this paper, it is important to note that operators with small OED exist in other bases as well [66, 67, 39, 34]. Therefore, studying the separation of other bases into equivalence classes may potentially yield more examples of simulable quantum dynamics.
Acknowledgements.
I. E. and O. L. are supported by Rosatom in the framework of the Roadmap for Quantum computing (Contract No. 868-1.3-15/15-2021 dated October 5,2021). Part of the work presented in paragraph “Disordered XY-spin chains” was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” under the grant N 22-1-2-55-1. T. B. is supported by the National Natural Science Foundation of China (62071301); NYU-ECNU Institute of Physics at NYU Shanghai; the Joint Physics Research Institute Challenge Grant; the Science and Technology Commission of Shanghai Municipality (19XD1423000,22ZR1444600); the NYU Shanghai Boost Fund; the China Foreign Experts Program (G2021013002L); the NYU Shanghai Major-Grants Seed Fund; Tamkeen under the NYU Abu Dhabi Research Institute grant CG008; and the SMEC Scientific Research Innovation Project (2023ZKZD55).